Determine the taxable amount of social security benefits for the following situations. If required, round your answers to the nearest dollar. If an amount is zero, enter "0".

a. Erwin and Eleanor are married and file a joint tax return. They have adjusted gross income of $42,600, no tax-exempt interest, and $14,910 of Social Security benefits. As a result, $ of the Social Security benefits are taxable.

b. Assume Erwin and Eleanor have adjusted gross income of $19,000, no tax-exempt interest, and $20,900 of Social Security benefits. As a result, $ of the Social Security benefits are taxable.

c. Assume Erwin and Eleanor have adjusted gross income of $119,500, no tax-exempt interest, and $17,925 of Social Security benefits. As a result, $ of the Social Security benefits are taxable.
Answer #1

For married filing jointly, the social security benefits are taxable to 50% for the combined income between $32,000 and $44,000.

For married filing jointly, the social security benefits are taxable to 85% for the combined income more than $44,000.

Part A

Taxable amount of social security benefits is lower of the following two:

Taxable amount of social security benefits = 14910*50% = $7455

Taxable amount of social security benefits = 42600+(14910*50%)-32000 = $18055

Therefore,

Taxable amount of social security benefits = $7455

Part B

Taxable amount of social security benefits is lower of the following two:

Taxable amount of social security benefits = 20900*50% = $10450

Taxable amount of social security benefits = 19000+(20900*50%)-32000 = -2550 = $0

Therefore,

Taxable amount of social security benefits = $0

Part C

Taxable amount of social security benefits is lower of the following two:

Taxable amount of social security benefits = 17925*85% = $15236

Taxable amount of social security benefits = 119500+(17925*85%)-44000 = $90736

Therefore,

Taxable amount of social security benefits = $15236